OpenStudy (anonymous):

Solve system of equations 5x+4y=6 -2x-3y=-1

OpenStudy (anonymous):

Help😭

OpenStudy (bradely):

5x+4y=6 ---- (1) -2x-3y=-1---- (2) 2(1)+5(2) 10x+8y=12 -10x-15y=-5 ---------- -7y=7 y=-1 5x+4(-1)=6 5x=10 x=2 source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Thank you @bradely